40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 836     Entry time: Thu Aug 14 19:08:14 2008
Author: Yoichi 
Type: Configuration 
Category: SUS 
Subject: Free swing measurement going on 
I started free swinging spectra measurement of all the suspensions now Aug 14 19:05 (PDT).
The watch dogs are all shutdown. Please do not turn them back on until tomorrow morning.
ELOG V3.1.3-